# Semantic Residual Prompts for Continual Learning

### v1

@@ 4.2 Ablation studies
STAR-Prompt $\mathbf{89.16}$$\pm$$0.38$ $\mathbf{86.50}$$\pm$$0.52$ $\mathbf{94.24}$$\pm$$0.09$ $\mathbf{67.39}$$\pm$$3.08$
@@
Classify with first-level keys $\boldsymbol{w}_{c}$ $87.35$$\pm$$0.12$ $85.99$$\pm$$0.38$ $90.52$$\pm$$1.03$ $58.41$$\pm$$3.40$
w/o first-level prompts $77.00$$\pm$$0.94$ $79.13$$\pm$$0.96$ $89.31$$\pm$$2.74$ $62.52$$\pm$$4.93$
@@
Prefix Tuning (no residuals) $84.19$$\pm$$0.09$ $81.04$$\pm$$0.50$ $91.54$$\pm$$0.43$ $65.01$$\pm$$1.27$
w/o Generative Replay $87.60$$\pm$$0.24$ $85.02$$\pm$$0.17$ $80.77$$\pm$$2.69$ $55.81$$\pm$$2.27$
w. Unimodal Generative Replay $88.48$$\pm$$0.21$ $85.50$$\pm$$0.17$ $90.96$$\pm$$0.90$ $57.91$$\pm$$0.79$
w/o Confidence Modulation $88.04$$\pm$$0.51$ $85.65$$\pm$$0.30$ $93.62$$\pm$$0.42$ $66.55$$\pm$$1.41$
Tab. 3 dissects the impact of each part of STAR-Prompt. For space constraints, we report one dataset for each treated domain. The results for the remaining datasets substantially confirm the following outcomes and are reported in the suppl. materials.
Two-level prompting.  The first two comparisons serves to evaluate the contribution of the two-level prompting strategy. Specifically, “Classify with first-level keys $\boldsymbol{w}_{c}$ indicates an ablative approach that lacks the second prompting stage. To classify, similarly to CoOp and AttriCLIP, it uses the learned keys $\boldsymbol{w}_{c}$ as class-prototypes (the posteriors are computed as in Eq. 9). Tab. 3 shows that, despite being consistently outperformed by the full STAR-Prompt, this strategy can achieve competitive performance, suggesting that the learned keys are actually stable. However, when more plasticity is required (as for EuroSAT and ISIC), the gap w.r.t. STAR-Prompt increases.
In a similar vein, the row “w/o first-level prompts” stands for a variation of STAR-Prompt that avoids learning the fist-level prompts. These are replaced with the static hand-crafted CLIP textual prompts proposed for ImageNet [32]. We use the resulting textual embedding as class-prototype keys, learning only second-level prompts $Q_{c}$. In contrast with the first ablative version, this one performs better on aerial/medical domains, highlighting the significance of second-level prompts in terms of plasticity.
Semantic residuals.  We hence show the results corresponding to replacing our semantic residual (Eq. 7) with Prefix Tuning. Specifically, following [46], for each ViT layer we train $5$ key and $5$ value prompt tokens, concatenated to the keys and the values of the respective MSA layer. As reported in Tab. 3, there is a disparity compared to STAR-Prompt, underscoring the advantage of our additive prompting mechanism.
Generative Replay.  When removing the generative replay stage, we observe a significant drop, especially when the task deviates from the pre-train. Moreover, row w. Unimodal Generative Replay shows the results obtained when the number of Gaussians $M$ collapses to 1 (Sec. 3.5). Despite it is generally better than w/o Generative Replay, the results are largely inferior to the $M=5$ used for the full model.
Confidence modulation of the semantic residuals.  Eq. 5 injects the confidence of the CLIP encoders in the residuals. The row w/o Confidence Modulation shows that this is beneficial across the different datasets.
@@ 5 Conclusion
This paper presents STAR-Prompt, a prompting method for Continual Learning based on three main novelties. First, we strengthen the stability of the prompt selection mechanism using a foundation model and two levels of prompt tuning. Second, we replace standard prompt token concatenation with additive residuals, which transfer semantics into its MLP layers. Finally, we use a simple generative replay based on a multi-modal representation of the feature distributions. Each part of STAR-Prompt brings a significant contribution, leading it to outperform the CL state of the art.

### v3

@@ 4.2 Ablation studies
STAR-Prompt $\mathbf{89.83}$$\pm$$0.04$ $\mathbf{87.62}$$\pm$$0.20$ $\mathbf{93.70}$$\pm$$0.09$ $\mathbf{66.67}$$\pm$$1.45$
@@
Classify with first-level keys $\boldsymbol{w}_{c}$ $88.16$$\pm$$0.27$ $87.57$$\pm$$0.12$ $90.12$$\pm$$0.54$ $58.87$$\pm$$1.25$
w/o first-level prompts $88.97$$\pm$$0.52$ $79.88$$\pm$$0.46$ $86.25$$\pm$$5.32$ $58.68$$\pm$$5.57$
@@
Prefix Tuning (no residuals) $71.34$$\pm$$0.34$ $60.03$$\pm$$4.14$ $90.92$$\pm$$0.44$ $62.46$$\pm$$0.88$
w/o Generative Replay $88.55$$\pm$$0.06$ $87.17$$\pm$$0.07$ $83.78$$\pm$$1.82$ $50.76$$\pm$$3.36$
w. Unimodal Generative Replay $89.62$$\pm$$0.12$ $87.28$$\pm$$0.12$ $92.58$$\pm$$0.40$ $60.79$$\pm$$0.68$
w/o Confidence Modulation $88.73$$\pm$$0.16$ $87.29$$\pm$$0.13$ $93.68$$\pm$$0.38$ $63.53$$\pm$$0.75$
Tab. 3 dissects the impact of each part of STAR-Prompt. For space constraints, we report one dataset for each domain (see suppl. materials for the others).
Two-level prompting.  Through the first two comparisons, we assess the role of the two-level strategy. Specifically, “Classify with first-level keys $\boldsymbol{w}_{c}$ indicates an ablative approach that lacks the second stage. To classify, similarly to CoOp, it uses the learned keys $\boldsymbol{w}_{c}$ as class-prototypes (the posteriors are computed as in Eq. 9). Tab. 3 shows that this strategy achieves competitive performance, highlighting the stability of the learned keys. However, if more plasticity is required (as for EuroSAT and ISIC), the gap w.r.t. STAR-Prompt increases. Similarly, STAR-Prompt shows improvement over the row labeled “w/o first-level prompts”, which is a variation of STAR-Prompt that avoids learning fist-level prompts and replaces them with the static hand-crafted CLIP textual templates [33]. We use the resulting textual embeddings as class-prototype keys, learning only second-level prompts $Q_{c}$.
Semantic residuals.  Afterwards, we replace our semantic residual (Eq. 7) with Prefix Tuning: following [48], for each ViT layer we train $5$ key and $5$ value prompt tokens, concatenated to the keys and the values of the respective MSA. Still, the results in Tab. 3 highlight the advantage of our additive mechanism.
Generative Replay.  A significant drop occurs if the generative replay stage is removed, especially when tasks deviate from the pre-train. Moreover, row w. Unimodal Generative Replay shows the results obtained when using a single Gaussian instead of a MoG. Despite it is generally better than w/o Generative Replay, the results are largely inferior to the MoG used for the full model.
Confidence modulation of the semantic residuals.  Eq. 5 injects the confidence of the CLIP encoders into the residuals. The last row shows that Confidence Modulation provides a fair improvement across the different datasets.
@@ 5 Conclusion
We present STAR-Prompt, a prompting method for Continual Learning based on three novelties. First, we strengthen the stability of prompt selection using a foundation model and two levels of prompt tuning. Second, we replace standard prompt concatenation with additive residuals, which transfer semantics into MLP layers. Finally, we use a simple generative replay based on a multi-modal representation of the feature distributions. Each part of STAR-Prompt brings a significant contribution, leading it to outperform the state of the art.
Acknowledgements
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high performance computing resources and support. This paper has been supported from Italian Ministerial grant PRIN 2020 “LEGO.AI: LEarning the Geometry of knOwledge in AI systems”, n. 2020TA3K9N. Additionally, the research activities of Angelo Porrello have been partially supported by the Department of Engineering “Enzo Ferrari” through the program FAR_2023_DIP – CUP E93C23000280005.
Appendix 0.A Prompt-selection stability
We hereby extend the results outlined in Sec. 3.4 and present outcomes for additional datasets/domains. Figure A provides a visual snapshot at the conclusion of the final task. It is evident that STAR-Prompt exhibits the highest precision in selecting the appropriate prompts. Figure B supplements this analysis with a detailed focus on the initial task. At first glance, CODA-Prompt [41] appears to match the retrieval performance of our approach in Split EuroSAT. However, we ascribe it to a bias of CODA-Prompt towards the prompts learned during the first task, as also observed in Fig. A. These additional results reaffirm our belief that STAR-Prompt strikes the optimal balance between stability and flexibility throughout the training sequence.
Refer to caption Refer to caption
Figure A: Analysis of prompt selection stability for Split Cars-196, Split EuroSAT, Split CUB-200, and Split CropDiseases. We assess various models at the end of the last task and report results as confusion matrices. The $y$ axis indicates the task of the query sample, while the $x$ axis shows the task of the corresponding selected key.
Appendix 0.B Additional methodological details
In this section, we provide the details concerning ${\cal L}_{\operatorname{GR}}^{Q}$, introduced in Sec. 3.5, and which closely follows the computation of ${\cal L}_{\operatorname{GR}}^{P}$. Specifically, similarly to Eq. 14, ${\cal L}_{\operatorname{GR}}^{Q}$ is defined as:
${\cal L}_{\operatorname{GR}}^{Q}=-\frac{1}{nNt}\sum_{c=1}^{Nt}\sum_{i=1}^{n}\log p(y_{i}=c|\tilde{\boldsymbol{e}}_{L,i}^{\operatorname{CLS}}),$ (A)
where $p(y_{i}=c|\tilde{\boldsymbol{e}}_{L,i}^{\operatorname{CLS}})$ is the score relating each synthetic representation $\tilde{\boldsymbol{e}}_{L,i}^{\operatorname{CLS}}\sim\mathcal{H}_{c}$ with the corresponding ground-truth class $c$. We compute it as:
$p(y_{i}=c|\tilde{\boldsymbol{e}}_{L,i}^{\operatorname{CLS}})=g_{\theta_{t^{\prime}}|_{t^{\prime}\in\{1,...,t\}}}(\tilde{\boldsymbol{e}}_{L,i}^{\operatorname{CLS}}).$ (B)
In Eq. B, $g_{\theta_{t^{\prime}}|_{t^{\prime}\in\{1,...,t\}}}(\cdot)$ indicates that we use all the classification heads of $g(\cdot)$ corresponding to the $t$ tasks so far observed.
Refer to caption Refer to caption
Figure B: Prompt retrieval on the first task of different datasets.
Appendix 0.C On the computational demands of prompt-based approaches
We use a serial two-stage training approach. First, we learn prompts for the CLIP text encoder, freeze them, and then learn prompts for the ImageNet-pre-trained ViT. Since each phase is independent, the overall cost stems from each respecting phase.
In the first stage, similar to [52], a batch of images requires one forward pass through the frozen CLIP image encoder and another through the text encoder to compute class-level textual embeddings. Due to the very low number of trainable parameters, this stage converges in a few epochs and requires no gradient computation for the image encoder.
In the second stage, we freeze the CLIP text encoder prompts and focus on the ViT prompts. The class prototypes from the text encoder can be cached, eliminating the need for additional forward passes through the text encoder. Again, each batch involves two forward passes: one through the CLIP image encoder and another through the prompted ViT. This complexity of STAR-Prompt thus matches L2P, Dual-Prompt, and CODA-Prompt, which also use two forward passes but with the same backbone. Our approach uses two distinct backbones, requiring additional GPU memory for the CLIP visual encoder, which we find negligible compared to the performance gain.
Appendix 0.D Setting and implementation details
Experimental setting. We here provide any missing information about datasets and experimental settings:
  • Split Imagenet-R: $10$ tasks of $20$ classes each; $50$ training epochs.
  • Split CIFAR-100: $10$ tasks of $10$ classes each; $20$ training epochs.
  • Split Cars-196: $9$ tasks of $20$ classes each, plus the 10th task comprising the remaining 16 classes; $50$ training epochs.
  • Split CUB-200: $10$ tasks of $20$ classes each; $50$ training epochs.
  • Split EuroSAT: $5$ tasks of $2$ classes each; $5$ training epochs.
  • Split RESISC45: $9$ tasks of $5$ classes each; $30$ training epochs.
  • Split CropDiseases: $7$ tasks of $5$ classes each; $5$ training epochs. From the original dataset [16] of $38$ classes, we removed the classes with the lowest number of examples (“Potato healthy”, “Apple Cedar” and “Peach healthy”).
  • Split ISIC: $3$ tasks of $2$ classes each; $30$ training epochs. From the original dataset [8] we removed the most frequent class “Melanocytic nevus”.
  • Split ChestX: $2$ tasks of $3$ classes each. $30$ training epochs. From the original dataset [47], we took the images without overlapping diseases belonging to the classes “Cardiomegaly”, “Consolidation”, “Edema”, “Fibrosis”, “Pleural Thickening” and “Pneumothorax”.
Following [50], each experiment is repeated $3$ times fixing the seeds $1993$, $1996$, and $1997$.
On instance/batch-wise inference. To maintain consistency with the majority of CL studies, we conduct inference independently for each sample within a batch. This approach, termed instance-wise inference, is by far the most predominant in the literature [24, 35, 1, 50, 41].
Conversely, L2P and DualPrompt originally presented results using a batch-wise setup, where a single prompt is selected for all samples in a batch through majority voting. We reckon that this setup offers an unfair advantage, as it leverages the fact that samples are not shuffled during inference, hence the ground-truth labels of the examples within a mini-batch are typically the same. To ensure a fair comparison with other methods, we thus report L2P and DualPrompt results using the instance-wise setup, thereby eliminating any potential advantage these approaches might have had over other techniques.
Appendix 0.E Final Forgetting Metric
Tabs. A and B report the Final Forgetting metric [5] for our experiments. Tab. A lacks L2P, DualPrompt, and SLCA, since [50] does not report forgetting values for their experiments. The same applies to PromptFusion and to the experiments of AttriCLIP on Split CIFAR-100 (taken from [46]).
Moreover, we report in Tab. C the standard deviation of experiments in Tab. 2 of the main paper.
Table A: The Final Forgetting (the lower the better) - part 1
Model Imagenet-R CIFAR-100 Cars-196 CUB-200 Avg.
Fine-tune  (ViT-B/16) $83.62$ $89.95$ $85.55$ $90.69$ $87.45$
LwF  $79.68$ $87.23$ $76.59$ $83.81$ $81.83$
DER++ * † $38.88$ $19.28$ $24.42$ $19.16$ $25.44$
CODA-Prompt $4.06$ $5.56$ $7.36$ $\mathbf{5.65}$ $5.66$
AttriCLIP $6.9$- $18.8$ $33.35$ $19.68$
STAR-Prompt $\mathbf{3.92}$ $\mathbf{4.71}$ $\mathbf{6.14}$ $6.94$ $\mathbf{5.43}$
Table B: The Final Forgetting (the lower the better) - part 2
Model EuroSAT RESISC CropDis. ISIC ChestX Avg.
Fine-tune  (ViT-B/16) $98.11$ $95.23$ $93.07$ $94.31$ $66.76$ $78.23$
LwF  $92.88$ $94.63$ $90.57$ $95.06$ $69.01$ $77.98$
DER++ * † $8.02$ $53.52$ $8.57$ $45.61$ $61.61$ $40.84$
L2P $43.87$ $25.41$ $18.85$ $37.79$ $42.60$ $36.88$
DualPrompt $12.88$ $14.46$ $10.98$ $\mathbf{25.01}$ $31.35$ $26.27$
CODA-Prompt $16.75$ $15.05$ $10.39$ $22.97$ $\mathbf{10.49}$ $22.41$
AttriCLIP $39.13$ $32.16$ $62.56$ $74.72$ $48.57$ $51.43$
SLCA  $7.74$ $10.58$ $4.90$ $35.25$ $32.05$ $27.30$
STAR-Prompt $\mathbf{4.31}$ $\mathbf{5.25}$ $\mathbf{3.26}$ $26.36$ $30.75$ $\mathbf{14.61}$
Table C: The std dev. for experiments in Tab. 2.
Model EuroSAT RESISC CropDis. ISIC ChestX
Joint (STAR-Prompt) $\pm$$0.13$$\pm$$0.24$$\pm$$0.10$$\pm$$0.30$$\pm$$0.43$
Joint  (ViT-B/16) $\pm$$0.12$$\pm$$0.10$$\pm$$0.08$$\pm$$1.76$$\pm$$1.40$
Fine-tune  (ViT-B/16) $\pm$$0.13$$\pm$$2.02$$\pm$$0.39$$\pm$2.222.$22$$\pm$$0.64$
LwF  $\pm$$2.78$$\pm$$0.85$$\pm$$2.76$$\pm$$1.98$$\pm$$0.85$
GDumb * † $\pm$$1.49$$\pm$$0.54$$\pm$$1.35$$\pm$$3.64$$\pm$$2.26$
DER++ * † $\pm$$1.62$$\pm$$2.89$$\pm$$1.06$$\pm$$2.16$$\pm$$1.22$
L2P $\pm$$7.86$$\pm$$3.71$$\pm$$0.25$$\pm$$3.84$$\pm$$1.52$
DualPrompt $\pm$$4.94$$\pm$$3.92$$\pm$$2.68$$\pm$$1.07$$\pm$$0.10$
CODA-Prompt $\pm$$6.30$$\pm$$5.15$$\pm$$2.91$$\pm$$3.50$$\pm$$3.90$
AttriCLIP $\pm$$6.15$$\pm$$4.31$$\pm$$3.03$$\pm$$10.38$$\pm$$1.81$
SLCA  $\pm$$0.48$$\pm$$0.35$$\pm$$0.60$$\pm$$3.83$$\pm$$1.80$
STAR-Prompt $\pm$$0.15$$\pm$$0.54$$\pm$$0.60$$\pm$$0.62$$\pm$$2.63$
Appendix 0.F Hyperparameters
The hyperparameters of STAR-Prompt employed for each experiment are reported in Tabs. D and E.
Refs. to Algorithm Imagenet-R CIFAR-100 Cars-196 CUB-200
$E_{1}$ First stage – Line 1 $10$ $10$ $10$ $50$
$\lambda$ First stage – Line 4 $30$ $10$ $30$ $30$
$lr$ First stage – Line 11 $0.05$ $0.05$ $0.01$ $0.05$
$E_{2}$ Second stage – Line 14 $10$ $10$ $10$ $5$
$\lambda$ Second stage – Line 17 $10$ $5$ $10$ $50$
$lr$ Second stage – Line 23 $0.001$ $0.001$ $0.01$ $0.1$
$M$ Requirements $5$ $5$ $5$ $5$
Table D: Hyperparameters used for each dataset - part 1.
Refs. to Algorithm EuroSAT RESISC CropDis. ISIC ChestX
$E_{1}$ First stage – Line 1 $10$ $10$ $10$ $50$ $10$
$\lambda$ First stage – Line 4 $30$ $10$ $30$ $5$ $30$
$lr$ First stage – Line 11 $0.05$ $0.05$ $0.01$ $0.01$ $0.05$
$E_{2}$ Second stage – Line 14 $10$ $10$ $10$ $10$ $10$
$\lambda$ Second stage – Line 17 $5$ $5$ $5$ $10$ $5$
$lr$ Second stage – Line 23 $0.1$ $0.01$ $0.001$ $0.01$ $0.05$
$M$ Requirements $5$ $5$ $5$ $5$ $5$
Table E: Hyperparameters used for each dataset - part 2.
Appendix 0.G Number of Trainable Parameters
The number of trainable parameters varies a lot among the compared methods. As mentioned in Secs. 1 and 2, the two main adaptation strategies are fine-tuning the whole model on the training data of the target dataset(s) or parameter-efficient learning, which adapts the model with only a few parameters (e.g., the prompts). In Tab. F we use Split CIFAR-100 as reference scenario and report the number of trainable parameters, i.e., those parameters that are optimized during the incremental learning. For instance, SLCA fine-tunes the whole model, while the trainable parameters of STAR-Prompt are composed of: $\boldsymbol{p}_{c},Q_{c},A_{c}$ ($c\in{\cal Y}_{1},...,{\cal Y}_{T}$) and $\theta_{t}$ ($t\in\{1,...,T\}$).
Model #params (millions)
Joint, Fine-Tune, LwF $86$
GDumb, DER++, SLCA $86$
L2P $0.12$
DualPrompt $0.41$
CODA-Prompt $3.91$
AttriCLIP $0.0998$
PromptFusion $0.35$
STAR-Prompt $3.89$
Table F: The number of learnable parameters for each tested method. Note that all these methods use a ViT-B/16 as the main classification architecture (see Sec. 4).
Appendix 0.H Additional ablations
We herein report the ablative studies of Sec. 4.2 over the remaining datasets in Tab. G. These results confirm the conclusion already outlined in the main paper (see Sec. 4.2).
Model CIFAR-100 CUB-200 RESISC CropDis. ChestX
STAR-Prompt 90.12$\pm$$0.32$ 84.10$\pm$$0.28$ 92.28$\pm$$0.54$ 94.92$\pm$$0.60$ 41.85$\pm$$2.63$
Ablations on two-level prompting
Classify with first-level keys $\boldsymbol{w}_{c}$ $83.49$$\pm$$0.16$ $81.31$$\pm$$0.20$ $90.80$$\pm$$0.34$ $88.36$$\pm$$0.74$ $31.70$$\pm$$0.77$
w/o first-level prompts $87.67$$\pm$$0.37$ $81.34$$\pm$$0.17$ $85.85$$\pm$$0.69$ $89.92$$\pm$$1.76$ $37.27$$\pm$$3.96$
Other secondary ablations
Prefix Tuning (no residuals) $86.70$$\pm$$0.59$ $82.67$$\pm$$0.14$ $84.26$$\pm$$0.35$ $95.06$$\pm$$0.34$ $39.28$$\pm$$4.14$
w/o Generative Replay $88.72$$\pm$$0.41$ $82.21$$\pm$$0.39$ $88.2$$\pm$$0.79$ $88.82$$\pm$$1.25$ $37.66$$\pm$$2.12$
w. Unimodal Generative Replay $90.07$$\pm$$0.34$ $83.16$$\pm$$0.14$ $92.29$$\pm$$0.51$ $94.21$$\pm$$0.44$ $38.77$$\pm$$2.76$
w/o Confidence Modulation $89.92$$\pm$$0.18$ $83.74$$\pm$$0.30$ $92.05$$\pm$$0.44$ $93.97$$\pm$$0.21$ $39.28$$\pm$$3.17$
Table G: Ablative studies on STAR-Prompt (Final Avg. Acc. ± std dev).
Appendix 0.I Additional experimental analysis
Number of Gaussians. Tab. H shows the Final Average Accuracy of STAR-Prompt on Split Imagenet-R using different numbers of Gaussians ($M$) for each $\mathcal{H}_{c}$. The results indicate a substantial plateau of the performance when $M\geq 5$, which we hence set as default value in all our experiments.
Prefix tuning vs. semantic residuals. In Tab. I we extend the comparison between prefix tuning and semantic residuals reported in Tab. 3 of the main paper. Specifically, differently from Prefix Tuning in Tab. 3, in which we used 5 tokens for each key and value, in Tab. I we use one token only for each key and value. This way, the total number of parameters per class is comparable with $Q_{c}$. Tab. I shows that the Final Average Accuracy scores obtained using prefix tuning are largely inferior to using semantic residuals, confirming the results reported in Tab. 3. This shows that the gap between the two conditioning methods is not due to the number of prompt parameters. The results of prefix tuning in I are drastically inferior to those obtained when using the recipe suggested in [48] (5 tokens for each key and value) and adopted in Tab. 3.
$M$ Final Average Accuracy
$2$ $88.94$
$5$ $89.37$
$10$ $89.16$
$20$ $89.58$
Table H: Impact of the number of Gaussians.
Conditioning method Imagenet-R CIFAR-100 Cars-196 CUB-200
Prefix Tuning (one token) $71.45$ $85.90$ $52.66$ $80.12$
Semantic Residuals 89.16 90.16 86.50 85.24
Table I: Comparing Semantic Residuals with Prefix Tuning. In the latter, we use a single-key and a single-value prompt token.